OpenStudy (cp123):

What is 2nd root of 27 plus the 2nd root of 3 equal to?

OpenStudy (cp123):

\[\sqrt{27} + \sqrt{3}\]

OpenStudy (cp123):

isnt it \[\sqrt[4]{3}\] ?

OpenStudy (cp123):

I just need someone to check my answer lol

OpenStudy (tkhunny):

Why would square roots turn into 4th roots? Simplify. No guessing. \(\sqrt{27} = \sqrt{3^3} = \sqrt{3^2\cdot 3} = 3\cdot\sqrt{3}\)

OpenStudy (cp123):

oh so would it be \[4\sqrt{3}\] i thought you added them together??

OpenStudy (tkhunny):

You DID add them together AFTER they we "like terms".

OpenStudy (cp123):

im confused so is it \[3\sqrt{3}\]

OpenStudy (tkhunny):

Why did you change your answer after you already had it right? 3x + x = 4x -- Like Terms "x" 3 Dogs + Dog = 4 Dogs -- Like Terms "Dogs" \(3\sqrt{3} + \sqrt{3} = 4\sqrt{3}\) -- Like Terms "\(\sqrt{3}\)" 3y + x = 3y + x -- Not Like Terms

OpenStudy (cp123):

oh sorry I read your response as if I did it wrong. I meant my question as questioning for confirmation, miscommunication on my behalf sorry.

OpenStudy (tkhunny):

I never do confirmation. Building confidence is part of what we are doing, here.

OpenStudy (cp123):

ok well thanks for the help